# RADLADS: Rapid Attention Distillation to Linear Attention Decoders at Scale

### v3

@@ Abstract
We present Rapid Attention Distillation to Linear Attention
Decoders at Scale (RADLADS), a protocol for rapidly converting softmax attention transformers into linear attention decoder models, along with two new RWKV-variant architectures, and models converted from popular Qwen2.5 open source models in 7B, 32B, and 72B sizes. Our conversion process requires only 350-700M tokens, less than 0.005% of the token count used to train the original teacher models. Converting to our 72B linear attention model costs less than $2,000 USD at today’s prices, yet quality at inference remains close to the original transformer. These models achieve state-of-the-art downstream performance across a set of standard benchmarks for linear attention models of their size. We release all our models on HuggingFace under the Apache 2.0 license, with the exception of our 72B models which are also governed by the Qwen License Agreement. 111Training Code: https://github.com/recursal/RADLADS-paper Models:
https://huggingface.co/collections/recursal/radlads-6818ee69e99e729ba8a87102
@@ 1 Introduction
Linear attention transformer variants employing compressive state have begun to match or even exceed traditional softmax attention based transformers in quality across many metrics. (Peng et al., 2025) This is important for long sequences at inference time, as linear attention is computable in O(1) time per token instead of O(N) time for softmax transformers, and avoids the expensive memory bandwidth usage of a Key Value cache. (Arora et al., 2023) However, the cost of training usable large models at scale is cost prohibitive for all but the largest organizations. This is especially true for large language models, where SoTA results have often required training on over ten trillion tokens of data. (Qwen et al., 2025) (Grattafiori et al., 2024)
Relative scorea Name Tokens Required Stages Lambada MMLU Others SUPRA (Mercat et al., 2024) 100B 1 0.913 0.216 0.869 LOLCats/Hedgehogb (Zhang et al., 2024a) 40M 2 - -0.288 0.635 MOHAWK (Bick et al., 2024b) 3-5B 3 0.938 -0.047 0.924 Mamba in the Llama (Wang et al., 2024) 20B 2 0.608 0.519 0.757 DiJiangb (Chen et al., 2024) 40B 1 - 0.720 - ARWKV (Yueyu et al., 2025) 60M/830M 2/3 0.893 0.801 0.935 Llamba (Bick et al., 2025) 8-12B 3 c0.951 c0.837 c0.987 RADLADS (ours) 350-700M 3 0.983 0.924 1.013
@@
Details of the RADLADS conversion process, including stepwise instructions, specific hyperparameters, token counts, and dataset choice.
@@
Public release of several converted Qwen models: QRWKV6-7B-Base, QRWKV6-7B-Instruct, QRWKV7-7B-Instruct, QRWKV6-32B-Instruct, QRWKV6-72B-Instruct, and QRWKV6-QwQ-32B.
@@ 2 Background
There have been many attempts at efficiently converting pre-trained transformer models to pure linear attention or other recurrent compressive state architectures. Early work, like Gerstenberger et al. (2020), performs full model logit distillation to a freshly initialized student model, and requires a very long training cycle. T2R (Kasai et al., 2021) attempts to reduce this training time by keeping most of the original model intact. It swaps softmax attention for traditional linear attention with a learnable MLP feature map, and finetunes the resultant model on approximately 3.5 billion tokens of data. Other projects like DiJiang (Chen et al., 2024), XATL (Choi, 2024), and SUPRA (Mercat et al., 2024) combine these two techniques, but still require 100B tokens of training or more, and exhibit lacking performance on some popular benchmarks like MMLU.
@@
Other works such as Mamba in the Llama (Wang et al., 2024) have focused on a pipeline of progressive distillation followed by supervised fine-tuning (Kim & Rush, 2016) and direct preference optimization (Rafailov et al., 2024). Mamba in the Llama uses 20B tokens of training. However, it focuses on hybrid models rather than pure RNNs, and performs poorly when the softmax attention is removed completely.
@@
We shared our early code, techniques, and 32B model with another team, who worked with those and subsequently published ARWKV (Yueyu et al., 2025), which converts Qwen2.5-7B-Instruct to use a standard RWKV-7 sequence mixer. As we discuss later, specific choices for weight transfer, hyperparameters, dataset, and architecture matter significantly. This is visible in the significantly higher downstream performance of our own Qwen2.5-7B-Instruct conversion, as shown in Tables 1 and 3.
@@ 3 Design Choices
Training for the correct number of tokens and with good choices for learning rates was also very important, both for efficiency of training and for final downstream performance. More training was not always better. We arrived at our final hyperparameters after theorizing that most knowledge resides in the teacher model’s MLPs and embeddings. Therefore, we start at a high learning rate to rapidly align the attention hidden states, annealing to approximately the final learning rate that the teacher model saw during pretraining. We tried various learning schedules, and cosine worked best. We keep the learning rate fixed during steps 2 and 3. This is an attempt to avoid catastrophic forgetting or overly significantly changing the knowledge encoded in the original teacher MLPs.
@@ 4 Method
In RADLADS, our student model architecture is an exact copy of the teacher model architecture, with the exception of the attention blocks. These are replaced with a choice of purely recurrent attention-replacement time mixing blocks. In this paper we detail two such replacements, which are based on RWKV6-C2 and RWKV7, which we name RAD-RWKV6 (”RADFinch”) and RAD-RWKV7(”RADGoose”) respectively. Our student models retain the runtime repetition of keys and values from Grouped Query Attention (GQA) (Ainslie et al., 2023) in the teacher model, if present. Depending on the attention-replacement chosen, we optionally include rotary positional embeddings (RoPE) (Su et al., 2023) from the teacher model. In cases where we include RoPE, it is applied at exactly the same point as in the teacher model.
See Appendix B for detailed formulas for RAD-RWKV6 and RAD-RWKV7.
@@ 4.2 Step 1: Attention Hidden State Alignment
The sequence length of this step is usually 512. We find that 100M tokens is enough to converge to a low stable loss during step 1. We run this step with a cosine annealed learning rate beginning at 1e-3 and ending at 1e-5. The final learning rate should be similar to the final learning rate the teacher saw during pretraining.
@@ 4.3 Step 2: Knowledge Distillation (Modelwise Approximation of Logits)
Empirically, we find that the model resulting from this step reproduces most of the initial model performance across a variety of benchmarks.
The sequence length of this step is usually 512. We run this step with a flat learning rate of 1e-5. This should be similar to the final learning rate the teacher saw during pretraining.
We find that 250M tokens is generally sufficient, and that training more than around 500M tokens during step 2 does not improve english language benchmarks. It’s possible that less frequently encountered tokens still improve beyond this point. This number is somewhat dependent on the quality of the attention replacement RNN used. For example, RAD-RWKV7 reaches better alignment during step 1 than RAD-RWKV6, and can therefore tolerate shorter step 2 training runs.
@@ 4.4 Step 3: Context Length Extension
We run this step with the same flat learning rate of 1e-5 as step 2.
@@
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.913 0.216 0.722 0.911 0.930 0.937 0.846 Mamba2-Llama3.0-8B-Instruct 0.608 0.519 0.724 0.868 0.901 0.903 0.391 MOHAWK-Phi1.5-1.3B 0.938 -0.047 0.830 0.968 0.936 0.959 0.927 DiJiang-Llama2.0-7B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - 0.720 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.973 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.732 0.906 0.975 - Llamba-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 0.951 0.837 0.983 1.014 0.969 0.994 0.975 LOLCatsHedgehog-Llama3.0-8B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - -0.029 0.534 0.864 0.750 0.886 0.143 ARWKV-7B 0.893 0.801 0.906 0.969 0.934 0.993 0.874 Our RADLADS models: QRWKV6-7B-Instruct 0.970 0.871 1.043 0.998 0.974 1.030 1.007 QRWKV6-7B-Instruct RoPE 0.969 0.880 1.054 1.003 0.971 1.018 0.981 QRWKV7-7B-Instruct 0.982 0.924 1.026 1.004 0.979 1.013 1.041 QRWKV7-7B-Instruct-from72B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 1.016 0.893 1.071 1.033 0.959 1.026 1.139 QRWKV6-32B-Instruct 0.986 0.909 1.066 1.037 0.964 1.005 1.235 QRWKV6-QwQ-32B 0.991 0.899 1.025 0.993 0.982 1.020 1.135 QRWKV6-72B-Instruct 1.004 0.899 1.015 1.008 0.974 0.968 1.123
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.692 0.331 0.458 0.759 0.771 0.801 0.703 teacher: Mistral-7B-v0.1 0.758 0.624 0.538 0.809 0.810 0.821 0.740 Mamba2-Llama3.0-8B-Instruct 0.439 0.452 0.480 0.741 0.708 0.758 0.586 teacher: Llama3.0-8B-instruct 0.722 0.639 0.567 0.816 0.758 0.786 0.719 MOHAWK-Phi1.5-1.3B 0.501 0.242 0.441 0.740 0.602 0.755 0.717 teacher: Phi1.5-1.3B-base 0.534 0.418 0.480 0.756 0.626 0.766 0.734 DiJiang-Llama2.0-7B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.407 0.427 0.626 0.694 0.775 - teacher: Llama2.0-7B - 0.468 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.463 c0.764 0.740 0.782 - Llamba-8Ba 0.694 0.610 0.546 0.825 0.776 0.809 0.733 1st teacher: Llama3.1-8B-Instruct 0.730 0.680 0.551 0.817 0.793 0.811 0.739 2nd teacher: Llama3.1-70B-Instruct 0.757 0.823 0.624 0.867 0.847 0.837 0.787 LoLCATsHedgehog-Llama3.0-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.238 0.401 0.726 0.656 0.765 0.533 teacher: Llama3.0-8B 0.666 0.533 0.801 0.791 0.799 0.731 0.381 ARWKV-7B 0.622 0.624 0.522 0.797 0.768 0.792 0.687 QRWKV6-7B-Instruct 0.675 0.657 0.563 0.814 0.790 0.803 0.711 QRWKV6-7B-Instruct-RoPE 0.674 0.661 0.567 0.817 0.789 0.799 0.706 QRWKV7-7B-Instruct-RoPE 0.684 0.682 0.558 0.817 0.793 0.798 0.718 teacher: Qwen2.5-7B-Instruct 0.696 0.717 0.550 0.815 0.805 0.794 0.714 QRWKV7-7B-Instruct-from72B111Not Available: Neither the model weights nor the missing scores were published by the model authors. 0.707 0.667 0.572 0.833 0.782 0.801 0.739 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763 QRWKV6-32B-Instruct 0.742 0.766 0.609 0.843 0.830 0.812 0.782 teacher: Qwen2.5-32B-Instruct 0.752 0.818 0.587 0.822 0.852 0.810 0.729 QRWKV6-QwQ-32B 0.662 0.743 0.564 0.806 0.830 0.804 0.732 teacher: Qwen2.5-QwQ-32B 0.668 0.799 0.556 0.810 0.841 0.798 0.705 QRWKV6-72B-Instruct 0.754 0.775 0.638 0.865 0.857 0.825 0.796 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763
@@ 5 Language Modeling Performance
In Table 3 we compare pure RNN models output from various conversion methods across a set of standard benchmarks. These converted models were trained from variety of different teacher models and vary in terms of parameter count. Therefore, we compare them via relative score: the ratio of benchmark accuracy scores between each student and teacher model, first subtracting the random chance accuracy for each benchmark. Negative ratios indicate that a model performed worse than random chance. Note that on nearly every benchmark, the RADLADS ”QRWKV” models obtain higher score ratios than models converted via any other method. (”QRWKV” is our shorthand for ”Qwen with RWKV attention”) In Table 3 QRWKV6-72B-Instruct demonstrates a new state-of-the-art in downstream performance for a pure RNN language model.
@@
Next, in Tables 5 and 5 we show models converted to hybrid-attention variants, each containing some amount of softmax attention. Even when compared to these hybrid models that still employ softmax attention, the pure RNN QRWKV models show the highest MMLU relative scores, and achieve similar or better relative scores on other benchmarks.
@@
Name Attention mmlu arc_c arc_e hella piqa winog LoLCATs-Mistralv0.1-7B SWA 100% 0.706 1.038 1.014 0.995 0.981 1.000 LoLCATs-Llama3.1-8B SWA 100% 0.727 1.037 1.015 1.001 1.069 0.839 LoLCATs-Llama3.1-70B SWA 100% 0.794 0.997 0.963 0.994 0.970 0.801 Mamba2-Llama3.0-8B-Instruct 12.5% 0.663 0.797 0.898 0.948 0.937 0.611 Mamba2-Llama3.0-8B-Instruct 25% 0.738 0.867 0.907 1.037 1.004 0.676 Mamba2-Llama3.0-8B-Instruct 50% 0.789 1.047 0.951 1.072 1.099 0.982
@@
Name Attention mmlu arc_c arc_e hella piqa winog LoLCATs-Mistralv0.1-7B SWA 100% 0.514 0.549 0.817 0.807 0.815 0.740 teacher: Mistralv0.1-7B 100% 0.624 0.538 0.809 0.810 0.821 0.740 LoLCATs-Llama3.1-8B SWA 100% 0.549 0.544 0.824 0.791 0.810 0.697 teacher: Llama-3.1-8B 100% 0.661 0.534 0.815 0.790 0.790 0.735 LoLCATs-Llama3.1-70B SWA 100% 0.677 0.605 0.850 0.846 0.821 0.737 teacher: Llama-3.1-70B 100% 0.788 0.606 0.873 0.850 0.831 0.796 Mamba2-Llama3.0-8B-Instruct 12.5% 0.508 0.503 0.758 0.731 0.768 0.634 Mamba2-Llama3.0-8B-Instruct 25% 0.537 0.525 0.764 0.777 0.787 0.648 Mamba2-Llama3.0-8B-Instruct 50% 0.557 0.582 0.788 0.795 0.815 0.715 teacher: Llama3.0-8B-Instruct 100% 0.639 0.567 0.816 0.758 0.786 0.719

### v4

@@ Abstract
We present Rapid Attention Distillation to Linear Attention Decoders at Scale (RADLADS), a protocol for rapidly converting softmax attention transformers into linear attention decoder models, along with two new RWKV-variant architectures, and models converted from popular Qwen2.5 open source models in 7B, 32B, and 72B sizes. Our conversion process requires only 350-700M tokens, less than 0.005% of the token count used to train the original teacher models. Converting to our 72B linear attention model costs less than $2,000 USD at today’s prices, yet quality at inference remains close to the original transformer. These models achieve state-of-the-art downstream performance across a set of standard benchmarks for linear attention models of their size. We release all our code [here] and models [here] under the Apache 2.0 license, with the exception of our 72B models which are also governed by the Qwen License Agreement.
@@ 1 Introduction
Linear attention transformer variants employing compressive state have begun to match or even exceed traditional softmax attention based transformers in quality across many metrics. (Peng et al., 2025) This is important for long sequences at inference time, as linear attention is computable in O(1) time per token instead of O(N) time for softmax transformers, and avoids the expensive memory bandwidth usage of a Key Value cache (Arora et al., 2023). However, the cost of training usable large models at scale is prohibitively costly for all but the largest organizations. This is especially true for large language models, where SoTA results have often required training on over ten trillion tokens of data. (Qwen et al., 2025) (Grattafiori et al., 2024)
Tokens Relative scorea (%) Name required Stages Lambada MMLU Others SUPRA (Mercat et al., 2024) 100B 1 91.3 21.6 86.9 LoLCats/Hedgehogb (Zhang et al., 2024a) 40M 2 - -28.8 63.5 MOHAWK (Bick et al., 2024b) 3-5B 3 93.8 -4.7 92.4 Mamba in the Llama (Wang et al., 2024) 20B 2 60.8 51.9 75.7 DiJiangb (Chen et al., 2024) 40B 1 - 72.0 - ARWKV (Yueyu et al., 2025) 60M/830M 2/3 89.3 80.1 93.5 Llamba (Bick et al., 2025) 8-12B 3 c95.1 c83.7 c98.7 RADLADS (ours) 350-700M 3 98.3 92.4 101.3
@@
Details of the RADLADS distillation recipe, including stepwise instructions, specific hyperparameters, token counts, and dataset choice.
@@
Public release of several converted Qwen models: QRWKV6-7B-Base, QRWKV6-7B-Instruct, QRWKV7-7B-Instruct, QRWKV6-32B-Instruct, QRWKV6-72B-Instruct.
@@ 2 Background
There have been many attempts at efficiently converting pre-trained transformer models to pure linear attention or other recurrent compressive state architectures. Early work (Gerstenberger et al., 2020) performed full model logit distillation to a freshly initialized student model, and required a very long training cycle. T2R (Kasai et al., 2021) attempts to reduce this training time by keeping most of the original model intact. It swaps softmax attention for traditional linear attention with a learnable MLP feature map, and finetunes the resultant model on approximately 3.5 billion tokens of data. Other projects like DiJiang (Chen et al., 2024), XATL (Choi, 2024), and SUPRA (Mercat et al., 2024) combine these two techniques, but still require 100B tokens of training or more, and exhibit lacking performance on some popular benchmarks like MMLU.
@@
Other works such as Mamba in the Llama (Wang et al., 2024) have focused on a pipeline of progressive distillation followed by supervised fine-tuning (Kim and Rush, 2016) and direct preference optimization (Rafailov et al., 2024). Mamba in the Llama uses 20B tokens of training. However, it focuses on hybrid models rather than pure RNNs, and performs poorly when the softmax attention is removed completely.
@@
We shared our early code, techniques, and 32B model with another team, who worked with those and subsequently published ARWKV (Yueyu et al., 2025), which converts Qwen2.5-7B-Instruct to use a standard RWKV-7 sequence mixer. As we discuss later, specific choices for weight transfer, hyperparameters, dataset, and architecture matter significantly. This is visible in the significantly higher downstream performance of our own Qwen2.5-7B-Instruct conversion, as shown in Tables 1 and 5.
@@ 3 Design choices
Training for the correct number of tokens and with good choices for learning rates was also very important, both for efficiency of training and for final downstream performance. More training was not always better. We arrived at our final hyperparameters after theorizing that most knowledge resides in the teacher model’s MLPs and embeddings. Therefore, we start at a high learning rate to rapidly align the attention hidden states, annealing to approximately the final learning rate that the teacher model saw during pretraining. We tried various learning schedules, and cosine worked best. We keep the learning rate for the MLPs fixed during steps 2 and 3. This is an attempt to avoid catastrophic forgetting or overly significantly changing the knowledge encoded in the original teacher MLPs.
@@ 4 Method
In RADLADS, our student model architecture is an exact copy of the teacher model architecture, with the exception of the attention blocks. These are replaced with a choice of purely recurrent attention-replacement sequence mixing blocks. In this paper we detail two such replacements, which are based on RWKV6-C2 and RWKV7, which we name RAD-RWKV6 (”RADFinch”) and RAD-RWKV7(”RADGoose”) respectively. Our student models retain the runtime repetition of keys and values from Grouped Query Attention (GQA) (Ainslie et al., 2023) in the teacher model, if present. Depending on the attention-replacement chosen, we optionally include rotary positional embeddings (RoPE) (Su et al., 2023) from the teacher model. In cases where we include RoPE, it is applied at exactly the same point as in the teacher model.
See Appendix A for detailed formulas for RAD-RWKV6 and RAD-RWKV7.
@@ 4.2 Step 1: Attention hidden state alignment
The sequence length of this step is usually 512. We find that 100M tokens is enough to converge to a low stable loss during step 1. We run this step with a cosine annealed learning rate beginning at 1e-3 and ending at 1e-5 or 7e-6. The final learning rate should be similar to the final learning rate the teacher saw during pretraining.
@@ 4.3 Step 2: Knowledge distillation (modelwise approximation of logits)
Empirically, we find that the model resulting from this step reproduces most of the initial model performance across a variety of benchmarks, but suffers from lack of context usage.
The sequence length of this step in our original protocol was 512. We ran this step with a flat learning rate of 1e-5 or 7e-6. This should be similar to the final learning rate the teacher saw during pretraining.
4.4 Step 2a: Knowledge distillation with context-length extension
More recently, we found that using a longer sequence length such as 4096 during step 2 can largely avoid the need for step 3. This step 2a is an optional replacement for steps 2 and 3. A naive run with increased length would take much longer and require much more data. But by supporting separate learning rates for the sequence mixer versus the rest of the model, we were able to adjust hyperparameters to reduce batch size and maintain the same total distillation token count and approximate speed. We find that keeping the sequence mixer learning rate high (1e-4) during most of step 2a and annealing it rapidly near the end of training and then continuing at a low rate (1e-5 or 7e-6), the model is able to adapt to the longer context length during distillation.
@@ 4.5 Step 3: Context length extension
We run this step with the same flat learning rate of 1e-5 or 7e-6 as step 2.
@@
5 Training details and hyper-parameters
We use the DCLM (Li et al., 2024b) dataset for all three steps in the RADLADS process, with the following hyper-parameters and token counts. For more optimizer details see Appendix B.
Step Tokens seqlen batch size adam betas, eps 1 100M cosine 512 32 0.9, 0.95, 2 250M-700M none 512 96 0.9, 0.95, 2a 600M s-cos-s a 4096 32 0.9, 0.95, 3 100M none 16384 96 0.9, 0.95,
is the learning rate applied to most of the model, while is the learning rate applied to the replacement sequence mixer. a In step 2a, cosine decay is employed between 83.3% and 87.5% of the tokens trained, with the learning rate remaining stable outside of that region.
@@ 6 Language modeling performance
In Table 5 we compare pure RNN models output from various conversion methods across a set of standard benchmarks. These converted models were trained from a variety of different teacher models and vary in terms of parameter count. Therefore, we compare them via relative score: the ratio of benchmark accuracy scores between each student and teacher model, first subtracting the random chance accuracy for each benchmark. Negative ratios indicate that a model performed worse than random chance. Note that on nearly every benchmark, the RADLADS ”QRWKV” models obtain higher score ratios than models converted via any other method. (”QRWKV” is our shorthand for ”Qwen with RWKV attention”) In Table 5 QRWKV6-72B-Instruct demonstrates a new state-of-the-art in downstream performance for a pure RNN language model.
@@
Next, in Tables 7 and 7 we show models converted to hybrid-attention variants, each containing some amount of softmax attention. Even when compared to these hybrid models that still employ softmax attention, the pure RNN QRWKV models show the highest MMLU relative scores, and achieve similar or better relative scores on other benchmarks.
Shown in Tables 10 and 11, we also tested a subset of models on harder and long context tasks gsm8k(Cobbe et al., 2021), HumanEval(Chen et al., 2021), passkey, and RULER(Hsieh et al., 2024). These demonstrate some of the limitations of distilling transformers to RNNs, especially for complex long sequence tasks.
To test reasoning performance, we distilled a RAD-RWKV7 model from the reasoning model Qwen3-8B. In Table 12, we show the Minerva Math (Hendrycks et al., 2021b; Lewkowycz et al., 2022) benchmark, a superset of the popular MATH500 (Lightman et al., 2023) test. Of interest is the fact that the converted model does not appear to improve as much with additional output tokens as the teacher model. This may indicate a need for improved context usage in the converted model.
Using DCLM we saw repetitive looping behaviors during reasoning, but found these can be eliminated by training on a mixture of 90% DCLM and 10% OpenThoughts (Guha et al., 2025), leading to a reasonable chain of thought. It is possible that reasoning datasets must be much more aligned with the teacher model than the ones we are using.
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 91.3 21.6 72.2 91.1 93.0 93.7 84.6 Mamba2-Llama3.0-8B-Instruct 60.8 51.9 72.4 86.8 90.1 90.3 39.1 MOHAWK-Phi1.5-1.3B 93.8 -4.7 83.0 96.8 93.6 95.9 92.7 DiJiang-Llama2.0-7B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - 72.0 333The DiJiang paper incorrectly lists 40.3 and 56.1 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.97.3 333The DiJiang paper incorrectly lists 40.3 and 56.1 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.73.2 90.6 97.5 - Llamba-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 5. 95.1 83.7 98.3 101.4 96.9 99.4 97.5 LoLCatsHedgehog-Llama3.0-8B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - -2.9 53.4 86.4 75.0 88.6 14.3 ARWKV-7B 89.3 80.1 90.6 96.9 93.4 99.3 87.4 Our RADLADS models: QLinAtt-7B-Instruct 90.7 64.0 94.6 96.7 90.7 98.0 90.2 QRWKV6-7B-Instruct 97.0 87.1 104.3 99.8 97.4 103.0 100.7 QRWKV6-7B-Instruct-RoPE 96.9 88.0 105.4 100.3 97.1 101.8 98.1 QRWKV7-7B-Instruct-RoPE 98.2 92.4 102.6 100.4 97.9 101.3 104.1 QRWKV7-7B-Instruct-from72B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 5. 101.6 89.3 107.1 103.3 95.9 102.6 113.9 QRWKV6-32B-Instruct 98.6 90.9 106.6 103.7 96.4 100.5 123.5 QRWKV6-72B-Instruct 1.004 89.9 101.5 100.8 97.4 96.8 112.3
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 69.2 33.1 45.8 75.9 77.1 80.1 70.3 teacher: Mistral-7B-v0.1 75.8 62.4 53.8 80.9 81.0 82.1 74.0 Mamba2-Llama3.0-8B-Instruct 43.9 45.2 48.0 74.1 70.8 75.8 58.6 teacher: Llama3.0-8B-Instruct 72.2 63.9 56.7 81.6 75.8 78.6 71.9 MOHAWK-Phi1.5-1.3B 50.1 24.2 44.1 74.0 60.2 75.5 71.7 teacher: Phi1.5-1.3B-base 53.4 41.8 48.0 75.6 62.6 76.6 73.4 DiJiang-Llama2.0-7B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 5. - 40.7 42.7 62.6 69.4 77.5 - teacher: Llama2.0-7B - 46.8 333The DiJiang paper incorrectly lists 40.3 and 56.1 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.46.3 c76.4 74.0 78.2 - Llamba-8Ba 69.4 61.0 54.6 82.5 77.6 80.9 73.3 1st teacher: Llama3.1-8B-Instruct 73.0 68.0 55.1 81.7 79.3 81.1 73.9 2nd teacher: Llama3.1-70B-Instruct 75.7 82.3 62.4 86.7 84.7 83.7 78.7 LoLCATsHedgehog-Llama3.0-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 5. - 23.8 40.1 72.6 65.6 76.5 53.3 teacher: Llama3.0-8B 66.6 53.3 80.1 79.1 79.9 73.1 38.1 ARWKV-7B 62.2 62.4 52.2 79.7 76.8 79.2 68.7 QLinAtt-7B-Instruct 63.1 54.9 53.4 79.6 75.3 78.8 68.9 QRWKV6-7B-Instruct 67.5 65.7 56.3 81.4 79.0 80.3 71.1 QRWKV6-7B-Instruct-RoPE 67.4 66.1 56.7 81.7 78.9 79.9 70.6 QRWKV7-7B-Instruct-RoPE 68.4 68.2 55.8 81.7 79.3 79.8 71.8 teacher: Qwen2.5-7B-Instruct 69.6 71.7 55.0 81.5 80.5 79.4 71.4 QRWKV7-7B-Instruct-from72B111Not Available: Neither the model weights nor the missing scores were published by the model authors. 70.7 66.7 57.2 83.3 78.2 80.1 73.9 teacher: Qwen2.5-72B-Instruct 75.1 83.4 63.2 85.9 87.4 83.6 76.3 QRWKV6-32B-Instruct 74.2 76.6 60.9 84.3 83.0 81.2 78.2 teacher: Qwen2.5-32B-Instruct 75.2 81.8 58.7 82.2 85.2 81.0 72.9 QRWKV6-72B-Instruct 75.4 77.5 63.8 86.5 85.7 82.5 79.6 teacher: Qwen2.5-72B-Instruct 75.1 83.4 63.2 85.9 87.4 83.6 76.3
@@
Name Attention mmlu arc_c arc_e hella piqa winog LoLCATs-Mistralv0.1-7B SWA 100% 70.6 1.038 1.014 99.5 98.1 100.0 LoLCATs-Llama3.1-8B SWA 100% 72.7 103.7 101.5 100.1 106.9 83.9 LoLCATs-Llama3.1-70B SWA 100% 79.4 99.7 96.3 99.4 97.0 80.1 Mamba2-Llama3.0-8B-Instruct 12.5% 66.3 79.7 89.8 94.8 93.7 61.1 Mamba2-Llama3.0-8B-Instruct 25% 73.8 86.7 90.7 103.7 100.4 67.6 Mamba2-Llama3.0-8B-Instruct 50% 78.9 104.7 95.1 107.2 109.9 98.2
@@
Name Attention mmlu arc_c arc_e hella piqa winog LoLCATs-Mistralv0.1-7B SWA 100% 51.4 54.9 81.7 80.7 81.5 74.0 teacher: Mistral-7B-v0.1 100% 62.4 53.8 80.9 81.0 82.1 74.0 LoLCATs-Llama3.1-8B SWA 100% 54.9 54.4 82.4 79.1 81.0 69.7 teacher: Llama-3.1-8B 100% 66.1 53.4 81.5 79.0 79.0 73.5 LoLCATs-Llama3.1-70B SWA 100% 67.7 60.5 85.0 84.6 82.1 73.7 teacher: Llama-3.1-70B 100% 78.8 60.6 87.3 85.0 83.1 79.6 Mamba2-Llama3.0-8B-Instruct 12.5% 50.8 50.3 75.8 73.1 76.8 63.4 Mamba2-Llama3.0-8B-Instruct 25% 53.7 52.5 76.4 77.7 78.7 64.8 Mamba2-Llama3.0-8B-Instruct 50% 55.7 58.2 78.8 79.5 81.5 71.5 teacher: Llama3.0-8B-Instruct 100% 63.9 56.7 81.6 75.8 78.6 71.9